OpenStudy (valirium):

What is the expression in factored form? 2x^2 + 16x + 24

OpenStudy (valirium):

@Directrix If this is a simple problem, then imma feel dumb lol

Directrix (directrix):

First, factor the common factor 2 out of each term. 2x^2 + 16x + 24 = 2 ( ? +? + ?)

Directrix (directrix):

Let's do this term by term. 2 times what = 2* x^2

OpenStudy (valirium):

Umm x?

Directrix (directrix):

Correct. Next, 2 times what = 8x

OpenStudy (valirium):

4x? Or just 4?

Directrix (directrix):

Whoops. I meant to ask this: 2 times what = 16x

Directrix (directrix):

Because I messed that up, I'll answer it. 2 times what = 16x 8x

Directrix (directrix):

Your turn: 2 times what = 24

OpenStudy (valirium):

12

Directrix (directrix):

Correct. Putting those into symbols: 2x^2 + 16x + 24 = 2* ( x^2 + 8x + 12) But, there is more factoring to do because: x^2 + 8x + 12 will factor

Directrix (directrix):

Can you think of two numbers that multiply to 12 AND add to 8 ?

OpenStudy (valirium):

Umm would it be 6 and 2?

Directrix (directrix):

Correct. x^2 + 8x + 12 will factor x^2 + 6x + 2x + 12 Does that type expansion look familiar for factoring?

OpenStudy (valirium):

Yess? haha

Directrix (directrix):

x^2 + 6x + 2x + 12 = x ( x + 6 ) + 2 ( x + 6) = ? (x+6) is the common factor here. See if you can finish up and don't forget to put the 2 out in front of these two factors. Let me know if you can get a wrap, okay?

Directrix (directrix):

@Valirium Question on finishing up?

OpenStudy (valirium):

So the answer is 2(x+6)(x+2)?

Directrix (directrix):

Correct